# On the Identifying Power of Generalized Monotonicity

### v2

@@ 1 Introduction
In their analysis of a setting with a binary outcome, treatment, and instrument, Balke and Pearl (1993, 1997) establish that the monotonicity condition of Imbens and Angrist (1994) has no identifying power beyond instrument exogeneity for average potential outcomes and the average treatment effect (ATE). Here, by no identifying power beyond instrument exogeneity, we mean that adding the monotonicity condition of Imbens and Angrist (1994) to instrument exogeneity does not decrease the identified sets for those parameters whenever those restrictions are consistent with the distribution of the observable data. In this way, their results contrast with the analysis of Imbens and Angrist (1994), who showed that their monotonicity condition and instrument exogeneity permitted identification of the local average treatment effect (LATE). This paper studies the extent to which this phenomenon holds in the broader context of a multi-valued outcome, treatment, and instrument.
@@
Our main result shows that a generalization of the monotonicity condition of Imbens and Angrist (1994) to this richer setting also has no identifying power beyond instrument exogeneity for average potential outcomes and ATEs. We hereafter refer to this condition more succinctly as generalized monotonicity. We show further that this result remains true for any restriction that is in fact stronger than generalized monotonicity provided that these stronger restrictions do not restrict potential outcomes in a sense that we will make precise later. Our analysis thereby accommodates many examples of restrictions on potential treatments that have been considered previously in the literature. In particular, we show that encouragement designs, the types of monotonicity restrictions considered by Kline and Walters (2016), Kirkeboen et al. (2016), and Heckman and Pinto (2018), and certain additive random utility models, including some studied in Lee and Salanié (2023), all satisfy generalized monotonicity. We further establish through a series of examples that models that do not satisfy generalized monotonicity can in fact have identifying power for average potential outcomes and ATEs. In this way, our analysis sheds light on the types of restrictions required for help in identifying average potential outcomes and ATEs.
@@
In establishing our results, we derive the identified sets for average potential outcomes under any such restriction and instrument exogeneity while maintaining the assumption that these restrictions are consistent with the distribution of the observable data. Our derivations reveal that the form of the resulting identified sets parallels the form of those derived by Balke and Pearl (1993, 1997) for a binary outcome, treatment, and instrument. An implication of the form of the identified sets is that average potential outcomes and ATEs are only identified under an identification-at-infinity-type condition when imposing instrument exogeneity and any such restriction.
@@
Our paper differs from the closely related literature that, in the context of a binary outcome, treatment, and instrument, considers the identifying power of the monotonicity condition of Imbens and Angrist (1994) and instrument exogeneity for the distribution (as opposed to the average) of potential outcomes, or considers the identifying power of these conditions when combined with additional restrictions on potential outcomes. In particular, Kamat (2019) shows that the monotonicity condition of Imbens and Angrist (1994) does have identifying power beyond instrument exogeneity for the distribution of potential outcomes. Machado et al. (2019) show that the monotonicity condition of Imbens and Angrist (1994) does have additional identifying power for the ATE beyond instrument exogeneity if one additionally imposes an assumption that requires potential outcomes to vary monotonically with the treatment. Thus, the phenomenon we explore is sensitive to both the choice of parameter and to whether one imposes assumptions on potential outcomes.
@@ 2 Setup and Notation
Denote by $Y\in\mathcal{Y}$ a multi-valued outcome of interest, by $D\in\mathcal{D}$ a multi-valued endogenous regressor, and by $Z\in\mathcal{Z}$ a multi-valued instrumental variable. To rule out degenerate cases, we assume throughout that $2\leq|\mathcal{Y}|<\infty$, $2\leq|\mathcal{D}|<\infty$, and $2\leq|\mathcal{Z}|<\infty$. Further denote by $Y_{d}\in\mathcal{Y}$ the potential outcome if $D=d\in\mathcal{D}$ and by $D_{z}\in\mathcal{D}$ the potential treatment if $Z=z\in\mathcal{Z}$. As usual, we assume that
@@
Let $P$ denote the distribution of $(Y,D,Z)$ and $Q$ denote the distribution of $((Y_{d}:d\in\mathcal{D}),(D_{z}:z\in\mathcal{Z}),Z)$. We will also denote by $Q|_{Y}$ the corresponding marginal distribution of $(Y_{d}:d\in\mathcal{D})$ under $Q$. For $T$ defined by (1), we have
@@
and therefore $P=QT^{-1}$.
Below we will require that $Q\in\mathbf{Q}$, where $\mathbf{Q}$ is a class of distributions satisfying assumptions that we will specify. Different choices of $\mathbf{Q}$ represent different assumptions that we impose on the distribution of potential outcomes and potential treatments. In this sense, $\mathbf{Q}$ may be viewed as a model for potential outcomes and potential treatments. By analogy with the notation introduced above, we will define $\mathbf{Q}|_{Y}=\{Q|_{Y}:Q\in\mathbf{Q}\}$. We say that $\mathbf{Q}$ does not restrict potential outcomes if and only if $\mathbf{Q}|_{Y}$ equals the entire set of distributions on $\mathcal{Y}^{|\mathcal{D}|}$.
Given $P$ and a model $\mathbf{Q}$, we define the set of $Q\in\mathbf{Q}$ that can rationalize $P$ as
$\mathbf{Q}_{0}(P,\mathbf{Q})=\{Q\in\mathbf{Q}:P=QT^{-1}\}~{}.$
We say $\mathbf{Q}$ is consistent with $P$ if and only if $\mathbf{Q}_{0}(P,\mathbf{Q})\neq\emptyset$. For every model $\mathbf{Q}$ considered in this paper, every $Q\in\mathbf{Q}$ is assumed to satisfy:
@@
$((Y_{d}:d\in\mathcal{D}),(D_{z}:z\in\mathcal{Z}))\perp\!\!\!\perp Z$ under $Q$.
Under this assumption, $\mathbf{Q}_{0}(P,\mathbf{Q})$ can be simplified as follows. If $Q$ satisfies Assumption 2.1, then
$p_{yd|z}\equiv P\{Y=y,D=d|Z=z\}=Q\{Y_{d}=y,D_{z}=d|Z=z\}=Q\{Y_{d}=y,D_{z}=d\}~{}.$ (2)
@@
$\mathbf{Q}_{0}(P,\mathbf{Q})=\bigg{\{}Q\in\mathbf{Q}:~{}p_{yd|z}=Q\{Y_{d}=y,D_{z}=d\},~{}\forall(y,d,z)\in\mathcal{Y}\times\mathcal{D}\times\mathcal{Z}\bigg{\}}~{}.$
@@
$\Theta_{0}(P,\mathbf{Q})\equiv\{\theta(Q):Q\in\mathbf{Q}_{0}(P,\mathbf{Q})\}~{}.$
@@
In what follows, we refer to Assumption 2.2 as generalized monotonicity. It states that under $Q$, for each treatment status $d\in\mathcal{D}$, there exists a value (possibly depending on $d$ and $Q$) of the instrument $z^{\ast}\in\mathcal{Z}$ that maximally encourages all individuals to $d$. Here, by “maximally encourage”, we mean that if an individual does not choose $d$ when $Z=z^{\ast}$, then they never choose $d$ for any other value of $Z$. Equivalently, if an individual chooses $d$ when $Z$ is equal to any value other than $z^{\ast}$, then they have to choose $d$ when $Z=z^{\ast}$. When $\mathcal{D}=\mathcal{Z}=\{0,1\}$, Assumption 2.2 is equivalent to the monotonicity assumption of Imbens and Angrist (1994). In this sense, Assumption 2.2 is a generalization of the monotonicity assumption of Imbens and Angrist (1994) to multiple treatments.
We emphasize that Assumption 2.2 only requires, for each possible value of the treatment, that there exists a value of the instrument that maximally encourages that treatment; it does not require that the value of the instrument is unique. For a given distribution $Q$ and given treatment $d\in\mathcal{D}$, let $\mathcal{Z}^{*}(d,Q)$ denote the set of $z^{\ast}$ that satisfy (3). In this notation, Assumption 2.2 can be restated as $\mathcal{Z}^{*}(d,Q)\neq\emptyset$ for each $d\in\mathcal{D}$. In the statement of Assumption 2.2, $z^{\ast}(d,Q)$ is allowed to change across $Q$. The following lemma shows that $\mathcal{Z}^{*}(d,Q)$ is identified from $P$ and is hence the same for all $Q$ that rationalizes $P$ and satisfies Assumptions 2.1 and 2.2. In what follows, we will therefore write $\mathcal{Z}^{*}(d)$ and $z^{*}(d)$ whenever the given distribution $Q$ rationalizes $P$ and satisfies Assumptions 2.1 and 2.2. This result generalizes the corresponding result in Imbens and Angrist (1994).
@@
Suppose $Q$ satisfies Assumptions 2.1 and 2.2 and $P=QT^{-1}$. Then, $z\in\mathcal{Z}^{*}(d,Q)$ if and only if
@@
Below we prove the necessity of (4); sufficiency is established in the appendix. Note that, for any $d\in\mathcal{D}$, $z^{\prime}\in\mathcal{Z}$, and any $z\in\mathcal{Z}^{*}(d,Q)$,
@@ 3 Main Result
In order to describe our main result, we first introduce some further notation. Denote by $\mathbf{Q}_{1}^{\ast}$ the set of all distributions that satisfy Assumption 2.1 and by $\mathbf{Q}_{1,2}^{\ast}$ the set of all distributions that satisfy Assumptions 2.1 and 2.2. In terms of this notation, our main result can be stated as follows:
@@
Suppose $\mathbf{Q}\subseteq\mathbf{Q}_{1,2}^{\ast}$ and $\mathbf{Q}|_{Y}$ equals the entire set of distributions on $\mathcal{Y}^{|\mathcal{D}|}$. Further suppose $\mathbf{Q}_{0}(P,\mathbf{Q})\neq\emptyset$. Then, $\Theta_{0}(P,\mathbf{Q})=\Theta_{0}(P,\mathbf{Q}_{1,2}^{\ast})=\Theta_{0}(P,\mathbf{Q}_{1}^{\ast})$.
Theorem 3.1 describes the sense in which restrictions on potential treatments stronger than generalized monotonicity have no identifying power for average potential outcomes and ATEs provided that these stronger restrictions do not restrict potential outcomes. This result is established through Theorems 3.2 and 3.3 below. Theorem 3.2, developed in Section 3.1, characterizes $\Theta_{0}(P,\mathbf{Q})$ for any model $\mathbf{Q}$ that is stronger than Assumptions 2.1 and 2.2, i.e., $\mathbf{Q}\subseteq\mathbf{Q}_{1,2}^{\ast}$, and does not restrict potential outcomes, i.e., for which $\mathbf{Q}|_{Y}$ equals the set of all distributions on $\mathcal{Y}^{|\mathcal{D}|}$. The result shows, in particular, that $\Theta_{0}(P,\mathbf{Q})=\Theta_{0}(P,\mathbf{Q}_{1,2}^{\ast})$ for any such model $\mathbf{Q}$ whenever $\mathbf{Q}_{0}(P,\mathbf{Q})\neq\emptyset$. Remarkably, this result holds even if the model $\mathbf{Q}$ is strictly more restrictive than Assumptions 2.1 and 2.2 in the sense that $\mathbf{Q}\subsetneqq\mathbf{Q}_{1,2}^{\ast}$. On the other hand, Theorem 3.3, developed in Section 3.2, shows that $\Theta_{0}(P,\mathbf{Q}_{1,2}^{\ast})=\Theta_{0}(P,\mathbf{Q}_{1}^{\ast})$ whenever $\mathbf{Q}_{0}(P,\mathbf{Q}_{1,2}^{\ast})\neq\emptyset$. Together, these results immediately imply Theorem 3.1.
3.1 Identified Sets for $\mathbf{Q}\subseteq\mathbf{Q}_{1,2}^{\ast}$
For $d\in\mathcal{D}$ and $z\in\mathcal{Z}$, define $\beta_{d|z}=\mathbb{E}_{P}[Y\mathbbm{1}\{D=d\}\mid Z=z]$. In addition, define $y^{L}=\min(\mathcal{Y})$ and $y^{U}=\max(\mathcal{Y})$. The following theorem derives the identified set for $\theta(Q)$, relative to any model that assumes instrument exogeneity and generalized monotonicity but does not restrict potential outcomes. Note in particular that the assumptions allow for $\mathbf{Q}\subsetneqq\mathbf{Q}_{1,2}^{\ast}$, in which case the model assumes strictly more than instrument exogeneity and generalized monotonicity.

### v3

@@ 1 Introduction
In their analysis of a setting with a binary outcome, treatment, and instrument, Balke and Pearl (1993, 1997) establish that the monotonicity condition of Imbens and Angrist (1994) has no identifying power beyond instrument exogeneity for average potential outcomes and the average treatment effect (ATE). Here, by no identifying power beyond instrument exogeneity, we mean that adding the monotonicity condition of Imbens and Angrist (1994) to instrument exogeneity does not decrease the identified sets for those parameters whenever those restrictions are consistent with the distribution of the observable data. [^1] For settings with possibly non-binary outcomes, Kitagawa (2021) shows this phenomenon continues to hold for any parameter that is a function of the marginal distributions of potential outcomes. In this way, their results contrast with the analysis of Imbens and Angrist (1994), who showed that their monotonicity condition and instrument exogeneity permitted identification of the local average treatment effect (LATE). This paper studies the extent to which this phenomenon holds in the broader context of a multi-valued outcome, treatment, and instrument.
@@
We show that a generalization of the monotonicity condition of Imbens and Angrist (1994) to this richer setting also has no identifying power beyond instrument exogeneity for average potential outcomes and ATEs. We hereafter refer to this condition more succinctly as generalized monotonicity. We show further that this result remains true for any restriction that is in fact stronger than generalized monotonicity provided that these stronger restrictions do not restrict potential outcomes in a sense that we will make precise later. This feature of our results is remarkable because one might expect stronger restrictions, possibly with very complicated restrictions on potential treatments that are difficult to fully characterize, to reduce the identified set at least in some instances, but we show that this is not the case. Using this remarkable result, our analysis accommodates many examples of restrictions on potential treatments that have been previously considered in the literature. In particular, we show that encouragement designs, the types of monotonicity restrictions considered by Kline and Walters (2016), Kirkeboen et al. (2016), and Heckman and Pinto (2018), and certain additive random utility models, including some studied in Lee and Salanié (2023), all satisfy generalized monotonicity. We further establish through a series of examples that models that do not satisfy generalized monotonicity can in fact have identifying power for average potential outcomes and ATEs. In this way, our analysis sheds light on the types of restrictions required for help in identifying average potential outcomes and ATEs.
@@
In establishing our results, we derive the identified sets for average potential outcomes under any such restriction and instrument exogeneity while maintaining the assumption that these restrictions are consistent with the distribution of the observable data. Our derivations reveal that the form of the resulting identified sets parallels the form of those derived by Balke and Pearl (1993, 1997) for a binary outcome, treatment, and instrument. An implication of the form of the identified sets is that average potential outcomes and ATEs are only identified under an identification-at-infinity-type condition when imposing instrument exogeneity and any such restriction. In our analysis, we also derive the identified sets for average potential outcomes and ATEs when imposing instrument exogeneity alone whenever the distribution of the observable data is consistent with instrument exogeneity and generalized monotonicity. As we explain in Example 4.1, this consistency is necessarily satisfied, for example, in the context of a multi-arm randomized controlled trial with one-sided non-compliance when defining the instrument to be random assignment to a given treatment arm.
@@
Our paper differs from the closely related literature that, in the context of a binary outcome, treatment, and instrument, considers the identifying power of the monotonicity condition of Imbens and Angrist (1994) and instrument exogeneity for the distribution (as opposed to the average) of potential outcomes, or considers the identifying power of these conditions when combined with additional restrictions on potential outcomes. In particular, Kamat (2019) shows that the monotonicity condition of Imbens and Angrist (1994) does have identifying power beyond instrument exogeneity for the (joint) distribution of potential outcomes. Machado et al. (2019) show that the monotonicity condition of Imbens and Angrist (1994) does have additional identifying power for the ATE beyond instrument exogeneity if one additionally imposes an assumption that requires potential outcomes to vary monotonically with the treatment. Thus, the phenomenon we explore is sensitive to both the choice of parameter and to whether one imposes assumptions on potential outcomes.
@@ 2 Setup and Notation
Denote by $Y\in\mathcal{Y}$ a multi-valued outcome of interest, by $D\in\mathcal{D}$ a multi-valued endogenous regressor, and by $Z\in\mathcal{Z}$ a multi-valued instrumental variable. [^2] Our restriction to a multi-valued $Y$ facilitates exposition, but is not essential. At the expense of slightly more complicated arguments, we can accommodate more generally any real-valued $Y$. To rule out degenerate cases, we assume throughout that $2\leq|\mathcal{Y}|<\infty$, $2\leq|\mathcal{D}|<\infty$, and $2\leq|\mathcal{Z}|<\infty$. Further denote by $Y_{d}\in\mathcal{Y}$ the potential outcome if $D=d\in\mathcal{D}$ and by $D_{z}\in\mathcal{D}$ the potential treatment if $Z=z\in\mathcal{Z}$. We impose the usual consistency assumption,
@@
Let $P$ denote the distribution of $(Y,D,Z)$ and $Q$ denote the distribution of $((Y_{d}:d\in\mathcal{D}),(D_{z}:z\in\mathcal{Z}),Z)$. Note that (1) defines a mapping $T$ through
@@
and therefore $P=QT^{-1}$. In what follows, we will say that a given $Q$ rationalizes a given $P$ if $P=QT^{-1}$.
Below we will require that $Q\in\mathbf{Q}$, where $\mathbf{Q}$ is a class of distributions satisfying assumptions that we will specify. Different choices of $\mathbf{Q}$ represent different assumptions that we impose on the distribution of potential outcomes and potential treatments. In this sense, $\mathbf{Q}$ may be viewed as a model for potential outcomes and potential treatments.
Given $P$ and a model $\mathbf{Q}$, the set of $Q\in\mathbf{Q}$ that can rationalize $P$ is
$\mathbf{Q}_{0}(P,\mathbf{Q})=\{Q\in\mathbf{Q}:P=QT^{-1}\}~{},$
i.e., the the pre-image of $P$ under $T$. We say $\mathbf{Q}$ is consistent with $P$ if and only if $\mathbf{Q}_{0}(P,\mathbf{Q})\neq\emptyset$. For every model $\mathbf{Q}$ considered in this paper, every $Q\in\mathbf{Q}$ is assumed to satisfy:
@@
Under $Q$, $(Y_{d},D_{z})\perp\!\!\!\perp Z$ for all $d\in\mathcal{D},z\in\mathcal{Z}$.
See Richardson and Robins (2013) for an analysis of alternative exogeneity assumptions, and in particular how this exogeneity assumption avoids the “cross-world” restrictions of the stronger joint independence assumption $((Y_{d}:d\in\mathcal{D}),(D_{z}:z\in\mathcal{Z}))\perp\!\!\!\perp Z$ under $Q$. Under this assumption, $\mathbf{Q}_{0}(P,\mathbf{Q})$ can be simplified as follows. If $Q$ satisfies Assumption 2.1, then
$p_{yd|z}:=P\{Y=y,D=d\mid Z=z\}=Q\{Y_{d}=y,D_{z}=d\mid Z=z\}=Q\{Y_{d}=y,D_{z}=d\}~{}.$ (2)
@@
$\mathbf{Q}_{0}(P,\mathbf{Q})=\bigg{\{}Q\in\mathbf{Q}:~{}p_{yd|z}=Q\{Y_{d}=y,D_{z}=d\}\text{ for all }y\in\mathcal{Y},d\in\mathcal{D},z\in\mathcal{Z}\bigg{\}}~{}.$
@@
$\Theta_{0}(P,\mathbf{Q}):=\{\theta(Q):Q\in\mathbf{Q}_{0}(P,\mathbf{Q})\}~{}.$
@@
In what follows, we refer to Assumption 2.2 as generalized monotonicity. It states that under $Q$, for each treatment status $d\in\mathcal{D}$, there exists a value (possibly depending on $d$ and $Q$) of the instrument $z^{\ast}\in\mathcal{Z}$ that maximally encourages all individuals to $d$. Here, by “maximally encourage”, we mean that if an individual does not choose $d$ when $Z=z^{\ast}$, then they never choose $d$ for any other value of $Z$. Equivalently, if an individual chooses $d$ when $Z$ is equal to any value other than $z^{\ast}$, then they have to choose $d$ when $Z=z^{\ast}$. When $\mathcal{D}=\mathcal{Z}=\{0,1\}$, Assumption 2.2 is equivalent to the monotonicity assumption of Imbens and Angrist (1994).
We emphasize that Assumption 2.2 only requires, for each possible value of the treatment, that there exists a value of the instrument that maximally encourages that treatment; it does not require that the value of the instrument is unique. For a given distribution $Q$ and given treatment $d\in\mathcal{D}$, let $\mathcal{Z}^{\ast}(d,Q)$ denote the set of $z^{\ast}$ that satisfy (3). In this notation, Assumption 2.2 can be restated as $\mathcal{Z}^{\ast}(d,Q)\neq\emptyset$ for each $d\in\mathcal{D}$. In the statement of Assumption 2.2, $z^{\ast}(d,Q)$ is allowed to change across $Q$. The following lemma shows that $\mathcal{Z}^{\ast}(d,Q)$ is identified from $P$ and is hence the same for all $Q$ that rationalizes $P$ and satisfies Assumptions 2.1 and 2.2. In what follows, we will therefore write $\mathcal{Z}^{\ast}(d)$ and $z^{\ast}(d)$ whenever the given distribution $Q$ rationalizes $P$ and satisfies Assumptions 2.1 and 2.2. This result generalizes the corresponding result in Imbens and Angrist (1994).
@@
Suppose $Q$ satisfies Assumptions 2.1 and 2.2 and $P=QT^{-1}$. Then, $z\in\mathcal{Z}^{\ast}(d,Q)$ if and only if
@@
Below we prove the necessity of (4); sufficiency is established in the appendix. Note that, for any $d\in\mathcal{D}$, $z^{\prime}\in\mathcal{Z}$, and any $z\in\mathcal{Z}^{\ast}(d,Q)$,
@@ 3 Main Result
In order to describe our main result, we first introduce some further notation. Denote by $\mathbf{Q}_{E}^{\ast}$ (where $E$ stands for exogeneity) the set of all distributions that satisfy Assumption 2.1 and by $\mathbf{Q}_{E,M}^{\ast}$ (where $M$ stands for generalized monotonicity) the set of all distributions that satisfy Assumptions 2.1 and 2.2. Denote by $Q|_{Y}$ the marginal distribution of $(Y_{d}:d\in\mathcal{D})$ under $Q$ and define $\mathbf{Q}|_{Y}=\{Q|_{Y}:Q\in\mathbf{Q}\}$. We will further require that the model does not restrict potential outcomes in the following sense:
Assumption 3.1 (Unrestricted Potential Outcomes).
$\mathbf{Q}|_{Y}$ equals the entire set of distributions on $\mathcal{Y}^{|\mathcal{D}|}$.
In terms of this notation, our main result can be stated as follows:
@@
Suppose $\mathbf{Q}\subseteq\mathbf{Q}_{E,M}^{\ast}$ and $\mathbf{Q}|_{Y}$ satisfies Assumption 3.1. Then, for any $P$ such that $\mathbf{Q}_{0}(P,\mathbf{Q})\neq\emptyset$, we have $\Theta_{0}(P,\mathbf{Q})=\Theta_{0}(P,\mathbf{Q}_{E,M}^{\ast})=\Theta_{0}(P,\mathbf{Q}_{E}^{\ast})$.
Theorem 3.1 describes the sense in which restrictions on potential treatments stronger than generalized monotonicity have no identifying power for average potential outcomes and ATEs provided that these stronger restrictions do not restrict potential outcomes. This result is established through Theorems 3.2 and 3.3 below. Theorem 3.2, developed in Section 3.1, characterizes $\Theta_{0}(P,\mathbf{Q})$ for any model $\mathbf{Q}$ that is stronger than Assumptions 2.1 and 2.2, i.e., $\mathbf{Q}\subseteq\mathbf{Q}_{E,M}^{\ast}$, and does not restrict potential outcomes in the sense of Assumption 3.1. The result shows, in particular, that $\Theta_{0}(P,\mathbf{Q})=\Theta_{0}(P,\mathbf{Q}_{E,M}^{\ast})$ for any such model $\mathbf{Q}$ whenever $\mathbf{Q}_{0}(P,\mathbf{Q})\neq\emptyset$. Remarkably, this result holds even if the model $\mathbf{Q}$ is strictly more restrictive than Assumptions 2.1 and 2.2 in the sense that $\mathbf{Q}\subsetneqq\mathbf{Q}_{E,M}^{\ast}$. On the other hand, Theorem 3.3, developed in Section 3.2, shows that $\Theta_{0}(P,\mathbf{Q}_{E,M}^{\ast})=\Theta_{0}(P,\mathbf{Q}_{E}^{\ast})$ whenever $\mathbf{Q}_{0}(P,\mathbf{Q}_{E,M}^{\ast})\neq\emptyset$. Together, these results immediately imply Theorem 3.1.
3.1 Identified Sets for $\mathbf{Q}\subseteq\mathbf{Q}_{E,M}^{\ast}$
For $d\in\mathcal{D}$ and $z\in\mathcal{Z}$, define $\beta_{d|z}=\mathbb{E}_{P}[Y\mathbbm{1}\{D=d\}\mid Z=z]$. In addition, define $y^{L}=\min(\mathcal{Y})$ and $y^{U}=\max(\mathcal{Y})$. The following theorem derives the identified set for $\theta(Q)$, relative to any model that assumes instrument exogeneity and generalized monotonicity but does not restrict potential outcomes. Note in particular that the assumptions allow for $\mathbf{Q}\subsetneqq\mathbf{Q}_{E,M}^{\ast}$, in which case the model assumes strictly more than instrument exogeneity and generalized monotonicity.